solve K= 1/2t(a-b) for a
Answer by nerdybill(7384)

solve K= 1/2t(a-b) for a
.
Start by multiplying both sides by 2:
2K = t(a-b)
Divide both sides by t:
2K/t = a-b
Finally add b to both sides:
2K/t + b = a